# Formal Degrees and Parabolic Induction: the Maximal Generic Case

### v2

@@ 6.1. The Formal Degree Conjecture
For simplicity, we shall still restrict ourselves to quasi-split groups. Thus $\mathbf{G}$ is still a quasi-split connected reductive group with $\mathbf{A}$ the split component of its center, and a fixed Borel subgroup $\mathbf{B}=\mathbf{T}\mathbf{U}$. Let $(\pi,V_{\pi})$ be a discrete series representation of $G$ with a fixed $G$-invariant inner product (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on $V_{\pi}$. Fix a Haar measure $\mu$ on $G/A$. Recall that the formal degree of $\pi$, with respect to $\mu$, is defined to be the unique positive real number $d(\pi)=d(\pi,\mu)\in\mathbb{R}^{\times}_{>0}$ such that
Clearly, $d(\pi)=d(\pi,\mu)$ is independent of the choice of (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) but depends on the choice of $\mu$, and $d(\pi,\lambda\mu)=\lambda^{-1}\cdot d(\pi,\mu)$.
@@
The explicit Plancherel formula for a general real reductive group has been completed by Harish-Chandra in 1970s, which was one of the highest and most fundamental achievements in representation theory of Lie groups. This was based on his explicit classification of the discrete series for all real semi-simple groups. However, such an explicit classification is inaccessible for general $p$-adic reductive groups due to the enormous arithmetic (Galois) complexity. In [HII08a], an explicit formula of $d(\pi,\mu)$ is conjectured in the framework of LLC. To state this conjecture in its precise form, we recall some desiderata of the refined LLC for quasi-split groups.
Let ${{}^{L}G}=\hat{G}\rtimes\Gamma_{K/F}$ be the L-group of $\mathbf{G}$ and $\hat{G}^{\natural}$ be the dual group of $\mathbf{G}/\mathbf{A}$. This is a subgroup of $\hat{G}$ containing $\hat{G}_{\mathrm{der}}$. Let $\varphi:\mathrm{WD}_{F}\rightarrow{{}^{L}G}$ be an L-parameter of $\mathbf{G}$ with the L-packet $\Pi_{\varphi}$. Let
$S_{\varphi}:=Z_{\hat{G}}(\varphi(\mathrm{WD}_{F})),~{}\mathcal{S}_{\varphi}:=\pi_{0}(S_{\varphi}/Z(\hat{G})^{\Gamma}),$
$S_{\varphi}^{\natural}:=Z_{\hat{G}^{\natural}}(\varphi(\mathrm{WD}_{F})),~{}\mathcal{S}_{\varphi}^{\natural}:=\pi_{0}(S_{\varphi}^{\natural}).$
Here $\mathcal{S}_{\varphi}$ is the “usual” component group used in the refined LLC for quasi-split groups, and $\mathcal{S}_{\varphi}^{\natural}$ is the component group used in the formal degree conjecture. Denote by $\mathrm{Irr}(H)$ the set of irreducible complex representations of a finite group $H$.
Conjecture 6.1 (Refined LLC for Quasi-Split Groups).
@@
    For each L-parameter $\varphi$, there should exist a bijection $\Pi_{\varphi}(\mathbf{G})\rightarrow\mathrm{Irr}(\mathcal{S}_{\varphi})$.
@@
    For tempered $\varphi$, there should exist a bijection $\iota_{\mathfrak{w}}:\Pi_{\varphi}(\mathbf{G})\rightarrow\mathrm{Irr}(\mathcal{S}_{\varphi})$ (depending on $\mathfrak{w}$), which maps the unique $\mathfrak{w}$-generic representation to the trivial representation of $\mathcal{S}_{\varphi}$.
@@
$\Pi_{\varphi}\rightarrow\mathrm{Irr}(\mathcal{S}_{\varphi}),~{}\pi^{\prime}\mapsto\rho_{\pi^{\prime}}$
as explained above. Recall that we’ve fixed a nontrivial character $\psi:F\rightarrow\mathbb{C}^{1}$. Let $\mu:=\mu_{2,\mathbf{G}}$ be the Haar measure on $G/A$ as explained at the beginning of §3, which only depends on the choice of $\psi$. Denote by $\mathrm{Ad}$ the adjoint representation of ${{}^{L}G}$ on $\mathrm{Lie}(\hat{G})/\mathrm{Lie}(Z(\hat{G})^{\Gamma})$. Then the formal degree conjecture reads:
Conjecture 6.2 ([HII08a][HII08]).
@@
  • Inner forms of $\mathrm{GL}(n)$: Silberger-Zink, cf. [HII08a, §4];
  • Unitary groups $\mathrm{U}(n)$: Beuzart-Plessis [Beu21] and Morimoto [Mor22];
  • Odd orthogonal groups $\mathrm{SO}(2n+1)$ and the metaplectic group $\mathrm{Mp}(2n)$: Ichino-Lapid-Mao [ILM17];
  • Kaletha’s regular and non-singular supercuspidals of tame groups: Schwein [Sch] and Ohara [Oha23];
@@
There was also an announcement by Beuzart-Plessis of a uniform proof for all classical groups via twisted endoscopy [Ca].
@@ 6.2. The Main Theorem
where this time the simple pole appears in the term $\gamma(s+js_{0},\sigma,r_{i},\bar{\psi})=\gamma(s+1,\sigma,r_{i},\bar{\psi})$.
@@
We close this section with some comments and examples. One may ask how to think about the quotient $\frac{m}{\langle\chi,\alpha^{\vee}\rangle}$ of some structure constants appearing in the above result. We want to remark that this quotient might come from the two different versions of component groups.
Recall that $\hat{G}^{\natural}$ is the dual group of $\mathbf{G}/\mathbf{A}$, identified as a subgroup of $\hat{G}$ containing $\hat{G}_{\mathrm{der}}$. We then have
$\hat{G}=\hat{G}^{\natural}\cdot Z(\hat{G})^{\Gamma}.$
Since $S_{\varphi}^{\natural}:=S_{\varphi}\cap\hat{G}^{\natural}$ and $S_{\varphi}\supset Z(\hat{G})^{\Gamma}$,
$S_{\varphi}=S_{\varphi}^{\natural}\cdot Z(\hat{G})^{\Gamma}.$
Then
$S_{\varphi}^{\natural}\twoheadrightarrow S_{\varphi}/Z(\hat{G})^{\Gamma}$
induces an isomorphism
$S_{\varphi}^{\natural}/(Z(\hat{G})^{\Gamma}\cap S_{\varphi}^{\natural})\simeq S_{\varphi}/Z(\hat{G})^{\Gamma}$
as complex Lie groups. By definition
$Z(\hat{G})^{\Gamma}\cap S_{\varphi}^{\natural}=Z(\hat{G})^{\Gamma}\cap S_{\varphi}\cap\hat{G}^{\natural}=Z(\hat{G}^{\natural})^{\Gamma},$
so we obtain
$S_{\varphi}^{\natural}/Z(\hat{G}^{\natural})^{\Gamma}\simeq S_{\varphi}/Z(\hat{G})^{\Gamma}.$
Since $\mathcal{S}_{\varphi}^{\natural}:=\pi_{0}(S_{\varphi}^{\natural})$, $\mathcal{S}_{\varphi}:=\pi_{0}(S_{\varphi}/Z(\hat{G})^{\Gamma})$, we get
$\frac{|\mathcal{S}_{\varphi}^{\natural}|}{|\pi_{0}(Z(\hat{G}^{\natural})^{\Gamma})|}=|\mathcal{S}_{\varphi}|,$
@@
(6.2) $\frac{|\mathcal{S}_{\varphi_{\pi}}^{\natural}|}{|\mathcal{S}_{\varphi_{\sigma}}^{\natural}|}=\frac{|\mathcal{S}_{\varphi_{\pi}}|}{|\mathcal{S}_{\varphi_{\sigma}}|}\cdot\frac{|\pi_{0}(Z(\hat{G}^{\natural})^{\Gamma})|}{|\pi_{0}(Z(\hat{M}^{\natural})^{\Gamma})|}.$
Here are some examples of split simple groups. We denote by $(a,b)$ the greatest common divisor of $a,b\in\mathbb{Z}$. The following table computes the quotient $\frac{m}{\langle\chi,\alpha^{\vee}\rangle}$.
$\mathbf{G}$ $\alpha$ $\mathbf{M}_{\alpha}$ $m$ $\chi\in X^{*}(\mathbf{M}_{\alpha})$ generator $\alpha^{\vee}$ $\frac{m}{\langle\chi,\alpha^{\vee}\rangle}$ $\mathrm{SL}_{n}$ $\varepsilon_{i}-\varepsilon_{i+1}$ $\mathrm{S}(\mathrm{GL}_{i}\times\mathrm{GL}_{n-i})$ $\frac{i(n-i)}{(n,i)}$ $(\det_{i}\boxtimes\mathbbm{1})\bmod\det_{n}$ $\varepsilon_{i}^{*}-\varepsilon_{i+1}^{*}$ $\frac{i(n-i)}{(n,i)}$ $\mathrm{PGL}_{n}$ $\varepsilon_{i}-\varepsilon_{i+1}$ $\mathrm{GL}_{i}\times\mathrm{GL}_{n-i}/\Delta\mathbb{G}_{m}$ $\frac{i(n-i)}{(n,i)}$ $\det_{i}^{\frac{n-i}{(n,i)}}\boxtimes\det_{n-i}^{\frac{-i}{(n,i)}}$ $\varepsilon_{i}^{*}-\varepsilon_{i+1}^{*}$ $\frac{i(n-i)}{n}$ $\mathrm{SO}_{2n+1}$ $\varepsilon_{n}$ $\mathrm{GL}_{n}$ $n$ $\det_{n}=\varepsilon_{1}+...+\varepsilon_{n}=2\tilde{\alpha}$ $2\varepsilon_{n}^{*}$ $\frac{n}{2}$ $\mathrm{SO}_{2n+1}$ $\varepsilon_{i}-\varepsilon_{i+1}$ $\mathrm{GL}_{i}\times\mathrm{SO}_{2(n-i)+1}$ $i$ $\det_{i}=\varepsilon_{1}+...+\varepsilon_{i}=\tilde{\alpha}$ $\varepsilon_{i}^{*}-\varepsilon_{i+1}^{*}$ $i$ $\mathrm{Sp}_{2n}$ $2\varepsilon_{n}$ $\mathrm{GL}_{n}$ $n$ $\det_{n}=\varepsilon_{1}+...+\varepsilon_{n}=\tilde{\alpha}$ $\varepsilon_{n}^{*}$ $n$ $\mathrm{Sp}_{2n}$ $\varepsilon_{i}-\varepsilon_{i+1}$ $\mathrm{GL}_{i}\times\mathrm{Sp}_{2(n-i)}$ $i$ $\det_{i}=\varepsilon_{1}+...+\varepsilon_{i}=\tilde{\alpha}$ $\varepsilon_{i}^{*}-\varepsilon_{i+1}^{*}$ $i$ $\mathrm{SO}_{2n}$ $\varepsilon_{n-1}+\varepsilon_{n}$ $\mathrm{GL}_{n}$ $n$ $\det_{n}=\varepsilon_{1}+...+\varepsilon_{n}=2\tilde{\alpha}$ $\varepsilon_{n-1}^{*}+\varepsilon_{n}^{*}$ $\frac{n}{2}$ $\mathrm{SO}_{2n}$ $\varepsilon_{i}-\varepsilon_{i+1}$ $\mathrm{GL}_{i}\times\mathrm{SO}_{2(n-i)}$ $i$ $\det_{i}=\varepsilon_{1}+...+\varepsilon_{i}=\tilde{\alpha}$ $\varepsilon_{i}^{\vee}-\varepsilon_{i+1}^{\vee}$ $i$ $\mathrm{G}_{2}$ long $\beta$ $\mathrm{GL}_{2}$ $2$ $\det=(3\alpha+2\beta)=\tilde{\beta}$ short $\beta^{\vee}$ $2$ $\mathrm{G}_{2}$ short $\alpha$ $\mathrm{GL}_{2}$ $2$ $\det=(2\alpha+\beta)=\tilde{\alpha}$ long $\alpha^{\vee}$ $2$
Remark.
In the above table, the case $\mathrm{SL}_{n},\mathrm{PGL}_{n}$ are a little subtler. The details are as follows.
  • For $\mathbf{G}=\mathrm{SL}_{n}$, $\alpha=\varepsilon_{i}-\varepsilon_{i+1}$, we identify
    $X^{*}(\mathbf{M}_{\alpha})=\frac{\mathbb{Z}(\varepsilon_{1}+...+\varepsilon_{i})\oplus\mathbb{Z}(\varepsilon_{i+1}+...+\varepsilon_{n})}{\mathbb{Z}(\varepsilon_{1}+...+\varepsilon_{n})},~{}X^{*}(\mathbf{A}_{\mathbf{M}_{\alpha}})=\frac{\mathbb{Z}e_{1}\oplus\mathbb{Z}e_{2}}{\mathbb{Z}(ie_{1}+(n-i)e_{2})},$
    and $X^{*}(\mathbf{M}_{\alpha})\rightarrow X^{*}(\mathbf{A}_{\mathbf{M}_{\alpha}})$ is given by
    $\overline{(\varepsilon_{1}+...+\varepsilon_{i})}\mapsto\overline{ie_{1}},~{}\overline{\varepsilon_{i+1}+...+\varepsilon_{n}}\mapsto\overline{(n-i)e_{2}},$
    thus of index $m=\frac{i(n-i)}{(n,i)}$.
  • For $\mathbf{G}=\mathrm{PGL}_{n}$, $\alpha=\varepsilon_{i}-\varepsilon_{i+1}$, we identify
    $X^{*}(\mathbf{M}_{\alpha})=\{a_{1}(\varepsilon_{1}+...+\varepsilon_{i})+a_{2}(\varepsilon_{i+1}+...+\varepsilon_{n}):ia_{1}+(n-i)a_{2}=0;~{}a_{1},a_{2}\in\mathbb{Z}\}$
    $X^{*}(\mathbf{A}_{\mathbf{M}_{\alpha}})=\{a_{1}e_{1}+a_{2}e_{2}:a_{1}+a_{2}=0;~{}a_{1},a_{2}\in\mathbb{Z}\}$
    and $X^{*}(\mathbf{M}_{\alpha})\rightarrow X^{*}(\mathbf{A}_{\mathbf{M}_{\alpha}})$ is given by
    $a_{1}(\varepsilon_{1}+...+\varepsilon_{i})+a_{2}(\varepsilon_{i+1}+...+\varepsilon_{n})\mapsto a_{1}ie_{1}+a_{2}(n-i)e_{2},$
    thus of index $m=\frac{i(n-i)}{(n,i)}$.
The next table computes $\frac{|\pi_{0}(Z(\hat{M}_{\alpha}^{\natural}))|}{|\pi_{0}(Z(\hat{G}^{\natural}))|}$ of these examples (where the Galois action is trivial since the groups are split).
$\hat{G}=\hat{G}^{\natural}$ $\alpha^{\vee}$ $\hat{M}_{\alpha}$ $\hat{M}_{\alpha}^{\natural}$ $\frac{|\pi_{0}(Z(\hat{M}_{\alpha}^{\natural}))|}{|\pi_{0}(Z(\hat{G}^{\natural}))|}$ $\mathrm{PGL}_{n}(\mathbb{C})$ $\varepsilon_{i}^{\vee}-\varepsilon_{i+1}^{\vee}$ $\mathrm{GL}_{i}\times\mathrm{GL}_{n-i}/\Delta\mathbb{G}_{m}$ $\mathrm{SL}_{i}\times\mathrm{SL}_{n-i}/\Delta\mu_{(n,i)}$ $\frac{i(n-i)}{(n,i)}$ $\mathrm{SL}_{n}(\mathbb{C})$ $\varepsilon_{i}^{\vee}-\varepsilon_{i+1}^{\vee}$ $\mathrm{S}(\mathrm{GL}_{i}\times\mathrm{GL}_{n-i})$ $\mathrm{SL}_{i}\times\mathrm{SL}_{n-i}$ $\frac{i(n-i)}{n}$ $\mathrm{Sp}_{2n}(\mathbb{C})$ $2\varepsilon_{n}^{\vee}$ $\mathrm{GL}_{n}$ $\mathrm{SL}_{n}$ $\frac{n}{2}$ $\mathrm{Sp}_{2n}(\mathbb{C})$ $\varepsilon_{i}^{\vee}-\varepsilon_{i+1}^{\vee}$ $\mathrm{GL}_{i}\times\mathrm{Sp}_{2(n-i)}$ $\mathrm{SL}_{i}\times\mathrm{Sp}_{2(n-i)}$ $i$ $\mathrm{SO}_{2n+1}(\mathbb{C})$ $\varepsilon_{n}^{\vee}$ $\mathrm{GL}_{n}$ $\mathrm{SL}_{n}$ $n$ $\mathrm{SO}_{2n+1}(\mathbb{C})$ $\varepsilon_{i}^{\vee}-\varepsilon_{i+1}^{\vee}$ $\mathrm{GL}_{i}\times\mathrm{SO}_{2(n-i)+1}$ $\mathrm{SL}_{i}\times\mathrm{SO}_{2(n-i)+1}$ $i$ $\mathrm{SO}_{2n}(\mathbb{C})$ $\varepsilon_{n-1}^{\vee}+\varepsilon_{n}^{\vee}$ $\mathrm{GL}_{n}$ $\mathrm{SL}_{n}$ $\frac{n}{2}$ $\mathrm{SO}_{2n}(\mathbb{C})$ $\varepsilon_{i}^{\vee}-\varepsilon_{i+1}^{\vee}$ $\mathrm{GL}_{i}\times\mathrm{SO}_{2(n-i)}$ $\mathrm{SL}_{i}\times\mathrm{SO}_{2(n-i)}$ $i$ $\mathrm{G}_{2}(\mathbb{C})$ $\beta^{\vee}$ $\mathrm{GL}_{2}$ $\mathrm{SL}_{2}$ $2$ $\mathrm{G}_{2}(\mathbb{C})$ $\alpha^{\vee}$ $\mathrm{GL}_{2}$ $\mathrm{SL}_{2}$ $2$
As we can see, for these basic examples (which are split simple of type $\mathrm{A}_{n-1},\mathrm{B}_{n},\mathrm{C}_{n},\mathrm{D}_{n},\mathrm{G}_{2}$), we have
(6.3) $\frac{m}{\langle\chi,\alpha^{\vee}\rangle}=\frac{|\pi_{0}(Z(\hat{G}^{\natural})^{\Gamma})|}{|\pi_{0}(Z(\hat{M}^{\natural})^{\Gamma})|}.$
Thus for these examples, by (6.2), the relation (6.1) is equivalent to
(6.4) $\frac{|\mathcal{S}_{\varphi_{\pi}}|}{|\mathcal{S}_{\varphi_{\sigma}}|}=j.$
But the author is not sure whether (6.3) holds in general under the setting of the main theorem.

### v3

@@ 6.1. The Formal Degree Conjecture
For simplicity, we shall still restrict ourselves to quasi-split groups. Thus $\mathbf{G}$ is still a quasi-split connected reductive group with $\mathbf{A}$ the split component of its center, and $\mathbf{B}=\mathbf{T}\mathbf{U}$ a fixed Borel subgroup. Let $(\pi,V_{\pi})$ be a discrete series representation of $G$ and we fix a Haar measure $\mu$. Recall that we denote by $d(\pi)=d(\pi,\mu)$ the formal degree of $\pi$ with respect to $\mu$ (cf. §3).
@@
The explicit Plancherel formula for a general real reductive group has been completed by Harish-Chandra in 1970s, which was one of the highest and most fundamental achievements in representation theory of Lie groups. This was based on his explicit classification of the discrete series for all real semi-simple groups. However, such an explicit classification is inaccessible for general $p$-adic reductive groups due to the enormous arithmetic (Galois) complexity. In [hiraga-ichino-ikeda], an explicit formula of $d(\pi,\mu)$ is conjectured in the framework of LLC. To state this conjecture in its precise form, we recall some desiderata of the refined LLC for quasi-split groups.
Let ${{}^{L}G}=\widehat{G}\rtimes\Gamma$ be the L-group of $\mathbf{G}$ and $\widehat{G}^{\natural}$ be the dual group of $\mathbf{G}/\mathbf{A}$. This is a subgroup of $\widehat{G}$ containing $\widehat{G}_{\mathrm{der}}$. Let $\varphi:\mathrm{WD}_{F}\rightarrow{{}^{L}G}$ be an L-parameter of $\mathbf{G}$ with the L-packet $\Pi_{\varphi}$. Let
$S_{\varphi}:=Z_{\widehat{G}}(\varphi(\mathrm{WD}_{F})),~{}\overline{S}_{\varphi}:=S_{\varphi}/Z(\widehat{G})^{\Gamma},$
$S_{\varphi}^{\natural}:=Z_{\widehat{G}^{\natural}}(\varphi(\mathrm{WD}_{F})),~{}\mathcal{S}_{\varphi}^{\natural}:=\pi_{0}(S_{\varphi}^{\natural}).$
The “usual” component group used in the refined LLC for quasi-split groups is $\pi_{0}(\overline{S}_{\varphi})$, and $\mathcal{S}_{\varphi}^{\natural}$ is the component group used in the formal degree conjecture. Here we should remark that since we only focus on quasi-split $\mathbf{G}$, the component group $\pi_{0}(\overline{S}_{\varphi})$ would suffice. For general connected reductive groups, let $\widehat{G}_{\mathrm{sc}}$ be the simply connected cover of the adjoint group $\widehat{G}_{\mathrm{ad}}:=\widehat{G}/Z(\widehat{G})$, then one has to use the pullback of $\overline{S}_{\varphi}$ in $\widehat{G}_{\mathrm{sc}}$ to define the component group $\mathcal{S}_{\varphi}$ (see [hiraga-ichino-ikeda, pp. 287]).
We denote by $\mathrm{Irr}(K)$ the set of irreducible complex representations of a finite group $K$.
Conjecture 6.1 (Refined LLC for $p$-adic Quasi-Split Groups).
@@
    For each L-parameter $\varphi$, there should exist a bijection $\Pi_{\varphi}(\mathbf{G})\rightarrow\mathrm{Irr}(\pi_{0}(\overline{S}_{\varphi}))$.
@@
    For tempered $\varphi$, there should exist a bijection $\iota_{\mathfrak{w}}:\Pi_{\varphi}(\mathbf{G})\rightarrow\mathrm{Irr}(\pi_{0}(\overline{S}_{\varphi}))$ (depending on $\mathfrak{w}$), which maps the unique $\mathfrak{w}$-generic representation to the trivial representation of $\pi_{0}(\overline{S}_{\varphi})$.
@@
$\Pi_{\varphi}\rightarrow\mathrm{Irr}(\pi_{0}(\overline{S}_{\varphi})),~{}\pi^{\prime}\mapsto\rho_{\pi^{\prime}}$
as explained above. Recall that we’ve fixed a nontrivial character $\psi:F\rightarrow\mathbb{C}^{1}$. Let $\mu:=\mu_{2,\mathbf{G}}$ be the Haar measure on $G/A$ as explained at the beginning of §3, which only depends on the choice of $\psi$. Denote by $\mathrm{Ad}$ the adjoint representation of ${{}^{L}G}$ on $\widehat{\mathfrak{g}}/Z(\widehat{\mathfrak{g}})^{\Gamma})$. Then the formal degree conjecture reads:
Conjecture 6.2 ([hiraga-ichino-ikeda][hiraga-ichino-ikeda-correction]).
@@
  • Inner forms of $\mathrm{GL}(n)$: Silberger-Zink, cf. [hiraga-ichino-ikeda, §4];
  • Unitary groups $\mathrm{U}(n)$: Beuzart-Plessis [bp-unitary] and Morimoto [morimoto];
  • Odd orthogonal groups $\mathrm{SO}(2n+1)$ and the metaplectic group $\mathrm{Mp}(2n)$: Ichino-Lapid-Mao [ichino-lapid-mao];
  • Kaletha’s regular and non-singular supercuspidals of tame groups: Schwein [schwein] and Ohara [ohara];
@@
There was also an announcement by Beuzart-Plessis of a uniform proof for all classical groups via twisted endoscopy [oberwolfach].
@@ 6.2. The Main Theorem
where this time the simple pole appears in the term $\gamma(s+js_{0},\sigma,r_{i},\psi)=\gamma(s+1,\sigma,r_{i},\psi)$.
@@
Here the structure constant $\langle\chi,\alpha^{\vee}\rangle/n$ comes from the modification of the component group in the formal degree conjecture. Let us just take a look at the simplest nontrivial example to see what happens.
Example.
Let $\mathbf{G}=\mathrm{GL}_{2n}$, $\mathbf{M}=\mathrm{GL}_{n}\times\mathrm{GL}_{n}$ be the usual standard Levi subgroup, and $\mathbf{P}=\mathbf{M}\mathbf{N}$ be the corresponding standard upper triangular parabolic subgroup of $\mathbf{G}$. By the Bernstein-Zelevinsky classification, pick an irreducible unitary supercuspidal $\tau$ of $\mathrm{GL}_{n}(F)$, write $\sigma:=\tau\boxtimes\tau$ as a representation of $M$, then there is a unique discrete series subrepresentation $\pi$ of
$i_{P}^{G}(\tau|\mathrm{det}_{n}|^{1/2}\boxtimes\tau|\mathrm{det}_{n}|^{-1/2}).$
We use the standard diagonal torus $\mathbf{T}$ and the standard coordinates $X^{*}(\mathbf{T})=\oplus_{i=1}^{2n}\mathbb{Z}e_{i}$. If we identify $X^{*}(\mathbf{M})$ with a subgroup of $X^{*}(\mathbf{T})$, then
$X^{*}(\mathbf{M})=\mathbb{Z}\cdot(e_{1}+\cdots+e_{n})\oplus\mathbb{Z}\cdot(e_{n+1}+\cdots+e_{2n})$
@@
$X^{*}(\mathbf{M})^{\mathbf{G}}=\mathbb{Z}\cdot[(e_{1}+\cdots+e_{n})-(e_{n+1}+\cdots+e_{2n})],$
with $m:=[X^{*}(\mathbf{A}_{\mathbf{M}})^{\mathbf{G}}:\mathrm{res}(X^{*}(\mathbf{M})^{\mathbf{G}})]=n$.
Of course the generator of $X^{*}(\mathbf{M})^{\mathbf{G}}$ is $\chi=(e_{1}+\cdots+e_{n})-(e_{n+1}+\cdots+e_{2n})$ and the coroot corresponding to our maximal parabolic $\mathbf{P}$ is $\alpha^{\vee}=e_{n}^{*}-e_{n+1}^{*}$. Now the fundamental weight $\tilde{\alpha}$ should be the element in $X^{*}(\mathbf{M})^{\mathbf{G}}\otimes\mathbb{R}=\mathfrak{a}_{\mathbf{M}}^{\mathbf{G}*}$ characterized by $\langle\tilde{\alpha},\alpha^{\vee}\rangle=1$, so that
$\tilde{\alpha}=\frac{1}{2}\cdot[(e_{1}+\cdots+e_{n})-(e_{n+1}+\cdots+e_{2n})]$
corresponding to $\det_{n}^{1/2}\boxtimes\det_{n}^{-1/2}$. So for this example $j=1$, and the structure constant above becomes
$\frac{\langle\chi,\alpha^{\vee}\rangle}{m}=\frac{2}{n}.$
This number comes from the modification of the “usual” component group. For $\mathbf{H}=\mathrm{GL}_{n}$, a discrete L-parameter is just an irreducible representation of $\varphi:\mathrm{WD}_{F}\rightarrow\mathrm{GL}_{n}(\mathbb{C})=\widehat{H}$, so that $Z_{\widehat{H}}(\varphi)=Z(\widehat{H})$ and $\pi_{0}(\overline{S}_{\varphi})=1$. But
$\widehat{H}^{\natural}:=(\mathbf{H}/\mathbf{A}_{\mathbf{H}})^{\wedge}=\mathrm{SL}_{n}(\mathbb{C})\subset\mathrm{GL}_{n}(\mathbb{C})=\widehat{H},$
which leads to $\mathcal{S}^{\natural}_{\varphi}=Z_{\widehat{H}^{\natural}}(\varphi)\simeq\mu_{n}$, so that in the above example
$\frac{|\mathcal{S}_{\varphi_{\pi}}^{\natural}|}{|\mathcal{S}_{\varphi_{\sigma}}^{\natural}|}=\frac{2n}{n\cdot n}=\frac{2}{n},$
which is exactly the constant $\langle\chi,\alpha^{\vee}\rangle/m$ above.